The solubility product (Ksp) or AgBr(s) is 2.0 x 〖10〗^(-13). Calculate the concentration of
soluble silver ion in a 0.05 M NaBr solution that was saturated with AgBr(s)?
1
Expert's answer
2018-05-16T07:13:59-0400
AgBr(s) → Ag+(aq) + Br-(aq)

Initial … 0 0.05M

Change … +1xM +xM

Equilibrium … xM x+0.05 M

Ksp = [Ag+][Br-] = x(x + 0.05) = x·0.05 = 2·10^-13

x = 4·10^-12

[Ag+] = 4·10^-12